what is 9 + r = -7
Found 2 solutions by uma, rcmcc:
Answer by uma(370)

9+r = -7
Adding -9 to both the sides we get,
r = -7-9
==> r = -16
Good Luck!!!

Answer by rcmcc(152)

9 + r = -7
solve for r
9 + r = -7
subtract 9 from both sides to isolate r
r=-7-9
r=-16